what will be approximate change in velocity by the tip of seconds hand of wall clock in 10s if length of the hand is 10cm

Grade:11

1 Answers

aman
6 years ago
v = wr [ where w is angular velocity and r is length of second hand]
and w = 2\pi/T [ where T is time period of second hand which is 60 mins so there for
w =\pi/30
so v =\pi/30 × 10/100 =  \pi/300 m/s
now \Deltav =\sqrt{}v+ v2 _2v2cos60   =   \pi/300 m/s